A physics instructor in an anteroom charges an electrostatic generator to a 25 micro columnb then it carries it in lecture room.The electric flux through lecture hall wall will be?
1
Expert's answer
2020-03-27T09:09:04-0400

The electric flux through lecture hall wall will be


"\Phi=\frac{q}{\epsilon_0}"

"\Phi=\frac{25\cdot 10^{-6}}{8.85\cdot 10^{-12}}=2.82\cdot 10^{6}\frac{F}{m}"